rosem13:

Sally Jones had a policy with a $250 deductible which paid 80% of her covered charges less deductible. She had medical expenses of $18,240.00. What is Sally's total cost?

jimthompson5910:

Here's a similar example: The problem is the same, but the numbers are different Question: Sally Jones had a policy with $300 deductible and paid 75% of her covered charges less deductible. She had medical expenses of $19,000. What is Sally's total cost?


Solution: The deductible is the amount Sally has to pay before insurance kicks in. Let's say her medical expenses were smaller than the deductible. This would mean she pays entirely for the medical bill and insurance pays nothing. Anything over $300 and insurance will kick in. Her bill was $19,000 and the deductible of $300 means 19,000-300 = 18,700 dollars is left for insurance to help out with. Insurance will cover 75% of this, which means, 75% of 18,700 = 0.75*18,700 = 14,025 dollars of the medical bill is covered by insurance. The remaining 18,700-14,025 = 4,675 dollars of the medical bill is not covered, so Sally will need to pay this herself. Note how 25% of 18,700 = 0.25*18700 = 4,675 which is the alternative way to figure out how much Sally must pay out of pocket (on top of the $300 deductible). To summarize: Sally has to deal with a $300 deductible plus the amount not covered by insurance, which was $4,765 Sally's total cost is 300+4,765 = 5,065 dollars Answer: $5,065 This won't be the answer to your exact question because I changed the numbers. Follow these steps to get the answer you're after.